# On the classification of centralizers of ODOs:

### v1

@@ 1 Introduction
The centralizer $\mathcal{Z}(L)$ of a nonzero ordinary differential operator (ODO) $L$ is a commutative algebra. This result was first proved by I. Schur in 1904 [45] and is one of the characteristic features of ordinary differential operators, with coefficients in a commutative differential ring, with no nonzero nilpotent elements [14]. Schur’s result does not hold for other algebras such as ODOs with matrix coefficients [40] or the algebra of partial differential operators [20].
The classification of all commutative algebras of ODOs has been studied by many authors and in diverse context of motivations, including Burchnall and Chaundy [5], Gelfand and Dikii [18], Krichever [22], Mumford [35], Segal–Wilson [46], Verdier [48], Mulase [34] and more recently Zheglov [53]. For ODOs, whose coefficients are formal power series, a bijective correspondence was established between the following categories of objects in [34]: The category of commutative algebras $\mathcal{A}$ of ordinary differential operators with a monic element, (for an invertible function $\sigma$, the algebras $\mathcal{A}$ and $\sigma\mathcal{A}\sigma^{-1}$ are identified); The moduli of isomorphism classes of quintets consisting of an algebraic curve and on it a point, a vector bundle, a local covering and a local trivialization. This is achieved via a contravariant functor, generalization of the Krichever map, establishing an equivalence between the category of arbitrary vector bundles on algebraic curves and the category of infinite dimensional vector spaces corresponding to certain points of Grassmannians together with their stabilizers.
@@
An algorithm has been designed in [19], to compute a finite basis of $\mathcal{Z}(L)$ as a $\mathbf{C}[L]$-module and to compute families of ODOs of arbitrary order with non-trivial centralizer $\mathcal{Z}(L)\neq\mathbf{C}[L]$, based on K. Goodearls results [14] and the computation of the Gelfand-Dickey hierarchies implemented in [10]. We use these bases and differential resultants to approach the computation of the defining ideal of ${\rm Spec}(\mathcal{Z}(L))$. So far it has been achieved successfully in the case of centralizers of ODOs of order $3$ in [44]. In this paper, we present results to extend our methods to commutative algebras $\mathcal{A}$ of ODOs, containing two operators of coprime order. These are algebras of rank $1$, being the rank the greatest common divisor of all orders. In general it is a non-trivial open problem to determine if an algebra of commuting operators has rank $1$, and although in many occasions this can be detected computing the centralizer, in some cases this is undecidable [19]. ODOs with centralizers of rank $1$ are also called algebro-geometric [49].
@@
The differential Galois group of $L-\lambda$, for a generic $\lambda$, was defined in [2], in the case of smooth spectral curves, and the centralizer of $L$ was proved to have rank $1$ if the group is commutative. The ultimate goal of our research project is to develop the Picard-Vessiot theory of spectral problems $L(y)=\lambda y$ for algebro-geometric ODOs, and give algorithms to compute differential Galois groups. Together with J.J. Morales Ruiz and M.A. Zurro, the spectral Picard-Vessiot field of $L(y)=\lambda y$ was described for an algebro-geometric Schrödinger operators $L$. This problem is linked to a factorization problem over a new coefficient field $\Sigma(\Gamma_{\mathcal{A}})$ defined by the spectral curve $\Gamma_{\mathcal{A}}$ of the centralizer $\mathcal{A}=\mathcal{Z}(L)$.
@@
The paper is organized as follows. In Section 2 we review results of K. Goodearl in [14] to describe the basis of a centralizer $\mathcal{Z}(L)$ as a $\mathbf{C}[L]$-module. For operators of prime order, Theorem 2.2 describes the basis of the centralizer and some conclusions are derived on the notion of ${\rm rank}(\mathcal{A})$ of a commutative subalgebra of $\Sigma[\partial]$. In general it is not easy to decide if a centralizer is trivial $\mathcal{Z}(L)=\mathbf{C}[L]$, but for differential operators of prime order in the first Weyl algebra, we can give an easy criteria Corollary 2.3.1.
@@
Section 3 is devoted to the description of an isomorphism between $\mathcal{A}$ and the coordinate ring of the spectral curve $\Gamma_{\mathcal{A}}$ in Theorem 3.3. This construction allows to define the ideal $\texttt{BC}(\mathcal{A})$ of the spectral curve, which is the ideal of all Burchnall-Chaundy polynomials. It was proved that it can be computed using differential resultants, so far for planar and space spectral curves in theorems 3.4 and 3.5.
@@
where $\partial^{-1}$ is the inverse of $\partial$ in $R((\partial^{-1}))$, $\partial^{-1}\partial=\partial\partial^{-1}=1$. We denote by $\mathcal{D}=\Sigma[\partial]$ the non-commutative ring of linear differential operators and recall that it is a principal ideal domain. In fact $\mathcal{D}$ admits Euclidean division. Given  $L$$M$ in $\mathcal{D}$, if $\rm ord(M)\geq\rm ord(L)$ then $M=qL+r$ with $\rm ord(r)<\rm ord(L)$, $q,r\in\Sigma[\partial]$. Let us denote by $\rm gcrd(L,M)$ the greatest common (right) divisor of $L$ and $M$.
@@ 2 Basis of centralizers
in the nontrivial case of operators $L\in\Sigma[\partial]\backslash\mathbf{C}[\partial]$. In most cases $\mathcal{Z}(L)=\mathbf{C}[L]$, but we will study the special case where $\mathbf{C}[L]$ is strictly included in $\mathcal{Z}(L)$ and say that the centralizer is non-trivial.
We will study the structure of $\mathcal{Z}(L)$ as a $\mathbf{C}[L]$-module, as it will allow us to describe $\mathcal{Z}(L)$ by a finite basis. For this purpose we will review next [14, Theorem 1.2]. Let us assume that $\rm ord(L)=n$ and define the $\mathbf{C}[L]$-submodules
@@
If $\mathcal{Z}_{i}$ is non empty then there exists $A_{i}$ in $\mathcal{Z}_{i}$ of minimal order. Observe that $\mathcal{Z}_{0}=\mathbf{C}[L]$ and $A_{0}=1$. By [14, Theorem 1.2], the centralizer $\mathcal{Z}(L)$ is a $\mathbf{C}[L]$-module with basis
@@
The rank of a commutative set of differential operators in $\Sigma[\partial]$ is the greatest common divisor of their orders. In particular, the rank of a pair of differentials operators is the greatest common divisor of their orders, see [39] for a discussion on the rank of a set of differential operators. As a consequence of Theorem 2.2, the rank of the centralizer,
@@
Given a subalgebra $\mathcal{A}$ of $\mathcal{Z}(L)$ determined by a finite set of generators,
$\mathcal{A}=\mathbf{C}[G_{0},G_{1},\ldots,G_{t-1}]$
then
${\rm rank}(\mathcal{A})=\rm gcd(\rm ord(G_{i})\mid i=0,\ldots,t-1).$
@@
Let $\mathcal{A}=\mathbf{C}[\mathcal{G}]$ be a finitely generated commutative subalgebra of $\Sigma[\partial]$, with finite set of generators $\mathcal{G}$ in $\Sigma[\partial]\backslash\mathbf{C}[\partial]$. The following are equivalent:
@@
    There exist operators $P$ and $Q$ in $\mathcal{G}$ of coprime order.
@@ 2.1 Computing ODOs with non-trivial centralizer
Given a differential operator $L$ in $\Sigma[\partial]$ it is an open problem to determine algorithmically if $\mathcal{Z}(L)$ is non-trivial. In addition, the examples of known families of ODOs with non-trivial centralizer are very limited, although profound and inspiring theoretical studies about their nature have been acheived throughout the twentieth century [23], [35],[51], [34],[53].
A crucial reason is that the coefficients of $L$ are solutions of nonlinear differential equations. If we restrict the case of monic ODOs in normal form (the coefficient of $\partial^{n-1}$ is zero)
@@
In [19], we use Goodearl’s Theorem and our SageMath implementation of GD hierarchies and almost commuting operators in [10] to compute centralizers.
@@
Consider the order $4$ operator given in  [37, Example 22]. The centralizer of the next operator of order $4$
@@ 2.2 In the first Weyl algebra
with $U(x)$, $V(x)$ and $W(x)$ polynomials in $\mathbb{C}[x]$. We proved in [39], Section 6 that given $L_{4}$ irreducible and such that $\mathcal{Z}(L_{4})\not=\mathbb{C}[L_{4}]$, then
$\mathcal{Z}(L_{4})={\mathbb{C}}[L,B_{2}]$
@@ 3 Spectral curves
By a famous theorem of I. Schur in [45], stated for analytic coefficients, given a monic differential operator $L$ in $\Sigma[\partial]$, if $\rm ord(L)=n$ then the centralizer $\mathcal{Z}((L))$ of $L$ in the ring of pseudo-differential operators is determined by the unique monic $n$-th root $L^{1/n}$ of $L$. For coefficients in a differential ring $R$ this result was proved in [14, Theorem 3.1]
@@
As a consequence, $\mathcal{Z}((L))$ is a commutative ring. To review the long history of this theorem, see  [14, Sections 3 and 4]. By  [14, Lemma 1.1] $\mathcal{Z}(L)$ is a domain and by  [14, Corollary 4.4] it is a commutative domain since
@@
Moreover any subalgebra $\mathcal{A}$ of $\mathcal{Z}(L)$ is the affine ring of an algebraic curve ${\rm Spec}(\mathcal{A})$. In this section, we aim to describe the defining ideal of this abstract curve.
@@ 3.1 Defining ideals of spectral curves
We will next define the Burchnall–Chaundy ideal of a finitely generated commutative $\mathbf{C}$-algebra $\mathcal{A}$ in $\Sigma[\partial]$. Without lost of generality can assume that
@@
with $\rm ord(L)<\rm ord(G_{i})$ and $G_{i}\notin\mathbf{C}[L]$. Observe that $\mathcal{A}\in\mathcal{Z}(L)$.
@@
We will call the elements of the BC ideal BC-polynomials.
Since we are working with differential operators in an euclidean domain $\Sigma[\partial]$ the following lemma follows.
Lemma 3.2.
The ideal $\texttt{BC}(\mathcal{A})$ is a prime ideal in $\mathbf{C}[\lambda,\overline{\mu}]$.
Proof.
To start $\texttt{BC}(\mathcal{A})$ is a nonzero ideal by [14], Theorem 1.13. For details, since $\mathcal{Z}(L)$ is a finitely generated $\mathbf{C}[L]$-module there exists $p_{0},\ldots,p_{t}\in\mathbf{C}[\lambda]$ such that
$g(\lambda,\mu_{i})=p_{0}(\lambda)+p_{1}(\lambda)\mu_{i}+\cdots+p_{t-1}(\lambda)\mu_{i}^{t-1}+\mu_{i}^{t}\in\texttt{BC}(\mathcal{A}).$
Given $h_{1}$ and $h_{2}$ in $\mathbf{C}[\lambda,\overline{\mu}]$, let us assume that $h_{1}\cdot h_{2}\in\texttt{BC}(\mathcal{A})$. Observe that $e_{\mathcal{A}}(h_{1})$ and $e_{\mathcal{A}}(h_{2})$ are differential operators in the domain $\Sigma[\partial]$. Since $e_{\mathcal{A}}$ is a ring homomorphism
$0=e_{\mathcal{A}}(h_{1}\cdot h_{2})=e_{\mathcal{A}}(h_{1})\cdot e_{\mathcal{A}}(h_{2})$
implies $e_{\mathcal{A}}(h_{1})=0$ or $e_{\mathcal{A}}(h_{2})=0$. Thus $h_{1}\in\texttt{BC}(\mathcal{A})$ or $h_{2}\in\texttt{BC}(\mathcal{A})$, proving that $\texttt{BC}(\mathcal{A})$ is prime. ∎
The next result explicitly shows that $\mathcal{A}$ is the coordinate ring of an algebraic curve whose defining ideal is $\texttt{BC}(\mathcal{A})$.
Theorem 3.3.
Given the finitely generated commutative $\mathbf{C}$-algebra $\mathcal{A}$ in $\Sigma[\partial]$ then the next isomorphism holds
@@
Proof.
We will denote the monomials in $\mathbf{C}[\lambda,\overline{\mu}]$ by
$M_{\alpha}=\lambda^{\alpha_{0}}\mu_{1}^{\alpha_{1}}\cdot\ldots\cdot\mu_{t-1}^{\alpha_{t-1}},\alpha=(\alpha_{0},\alpha_{1},\ldots,\alpha_{t-1})\in\mathbb{N}^{t},$
whose image is $e_{\mathcal{A}}(M_{\alpha})=L^{\alpha_{0}}G_{1}^{\alpha_{1}}\cdot\ldots\cdot G_{t-1}^{\alpha_{t-1}}$. Thus the image of $e_{L}$ is
$\mathcal{A}=\mathbf{C}[L,G_{1},\ldots,G_{t-1}]=\left\{\sum_{\alpha\in\mathbb{N}^{t}}\sigma_{\alpha}e_{L}(M_{\alpha})\ \mid\ \sigma_{\alpha}\in\mathbf{C}\right\}.$
Recall that $\texttt{BC}(\mathcal{A})=\rm Ker(e_{\mathcal{A}})$ to obtain the isomorphism. ∎
Thus $\mathcal{A}$ is the affine ring of an algebraic curve whose defining ideal is $\texttt{BC}(\mathcal{A})$. The spectral curve of $\mathcal{A}$ is the irreducible algebraic variety of the prime ideal $\texttt{BC}(\mathcal{A})$, that is
@@
$\mathbf{C}[\Gamma_{\mathcal{A}}]=\frac{\mathbf{C}[\lambda,\overline{\mu}]}{\texttt{BC}(\mathcal{A})}.$

### v2

@@ 1 Introduction
The centralizer $\mathcal{Z}(L)$ of a nonzero ordinary differential operator (ODO) $L$, in the ring of differential operators, is a commutative algebra. This result was first proved by I. Schur in 1904 [54] and is one of the characteristic features of ordinary differential operators, with coefficients in a commutative differential ring, with no nonzero nilpotent elements [19]. Schur’s result does not hold for other algebras such as ODOs with matrix coefficients [49] or the algebra of partial differential operators [27], [64].
The classification of all commutative algebras of ODOs has been studied by many authors and in diverse context of motivations, including Burchnall and Chaundy [7], Gelfand and Dikii [24], Krichever [29], [30], Mumford [44], Segal–Wilson [55], Previato-Wilson [50], Verdier [58], Mulase [43] and more recently Zheglov [63]. For ODOs, whose coefficients are formal power series, a bijective correspondence was established between the following sets of objects in [43]: The set of equivalence classes of commutative algebras $\mathcal{A}$ of ordinary differential operators with a monic element, (for an invertible function $\sigma$, the algebras $\mathcal{A}$ and $\sigma\mathcal{A}\sigma^{-1}$ are identified); The set of equivalence classes of isomorphism classes of quintets consisting of an algebraic curve and on it a point, a vector bundle, a local covering and a local trivialization. This is achieved via a contravariant functor, generalization of the Krichever map, establishing an equivalence between the category of quintets and the category of infinite dimensional vector spaces corresponding to certain points of Grassmannians together with their stabilizers, the so called Schur pairs [63]. Mulase’s work [43] was the last one of a long research history, more details can be found in the introduction of [50] or [63].
@@
An algorithm has been designed in [25], to compute a finite basis of $\mathcal{Z}(L)$ as a $\mathbf{C}[L]$-module and to compute families of ODOs of arbitrary order with non-trivial centralizer $\mathcal{Z}(L)\neq\mathbf{C}[L]$, based on K. Goodearl’s results [19] and the computation of the Gelfand-Dickey hierarchies implemented in [26]. We use these bases and differential resultants to approach the computation of the defining ideal of ${\rm Spec}(\mathcal{Z}(L))$. So far it has been achieved successfully in the case of centralizers of ODOs of order $3$ in [53]. In this paper, we present results to extend our methods to commutative algebras $\mathcal{A}$ of ODOs, containing two operators of coprime order. These are algebras of rank $1$, being the rank the greatest common divisor of all orders. ODOs with centralizers of rank $1$ are also called algebro-geometric [59]. In general, it is non-trivial to determine effectively if an algebra of commuting operators has rank $1$. In many occasions this can be detected through an effective computation of the centralizer [25], but in some cases, the algorithm in [25] cannot guarantee to compute the whole centralizer but only a maximal subalgebra of rank $r$, being $r$ the rank of an initially known pair of commuting operators.
@@
The differential Galois group of $L-\lambda$ for a generic $\lambda$, was defined in [3] and the centralizer of $L$ was proved to have rank $1$ if the group is commutative. The ultimate goal of our research project is to develop the Picard-Vessiot theory of spectral problems $L(y)=\lambda y$ for algebro-geometric ODOs, and give algorithms to compute differential Galois groups. Together with J.J. Morales Ruiz and M.A. Zurro, the spectral Picard-Vessiot field of $L(y)=\lambda y$ was described for an algebro-geometric Schrödinger operators $L$ [42]. This problem is linked to a factorization problem over a new coefficient field $\Sigma(\Gamma_{\mathcal{A}})$ defined by the spectral curve $\Gamma_{\mathcal{A}}$ of the centralizer $\mathcal{A}=\mathcal{Z}(L)$, whose field of constants is not algebraically closed.
@@
The paper is organized as follows. In Section 2 we review results of K. Goodearl in [19] to describe the basis of a centralizer $\mathcal{Z}(L)$ as a $\mathbf{C}[L]$-module. For operators of prime order, Theorem 2.2 describes the basis of the centralizer and some conclusions are derived on the notion of ${\rm rank}(\mathcal{Z}(L))$. In general it is not easy to decide if a centralizer is trivial $\mathcal{Z}(L)=\mathbf{C}[L]$, but for differential operators of prime order in the first Weyl algebra, we can give an easy criteria Corollary 2.3.1.
@@
Section 3 is devoted to the description of an isomorphism between $\mathcal{A}$ and the coordinate ring of the spectral curve $\Gamma_{\mathcal{A}}$. This construction allows to define the ideal $\texttt{BC}(\mathcal{A})$ of the spectral curve, which is the ideal of all Burchnall-Chaundy polynomials. We review some classical results about the computation of $\texttt{BC}(\mathcal{A})$ using differential resultants in the case of planar spectral curves in Theorem 3.2 and recent results on the case of space spectral curves in Theorem 3.3.
@@
where $\partial^{-1}$ is the inverse of $\partial$ in $R((\partial^{-1}))$, $\partial^{-1}\partial=\partial\partial^{-1}=1$. We denote by $\mathcal{D}=\Sigma[\partial]$ the non-commutative ring of linear differential operators and recall that its ideals are principal. In fact $\mathcal{D}$ admits Euclidean division. Given  LLMM in $\mathcal{D}$, if $\rm ord(M)\geq\rm ord(L)$ then $M=qL+r$ with $\rm ord(r)<\rm ord(L)$, $q,r\in\Sigma[\partial]$. Let us denote by $\rm gcrd(L,M)$ the greatest common (right) divisor of $L$ and $M$.
@@ 2 Basis of centralizers
in the non-trivial case of operators $L\in\Sigma[\partial]\backslash\mathbf{C}[\partial]$. In most cases $\mathcal{Z}(L)=\mathbf{C}[L]$, but we will study the special case where $\mathbf{C}[L]$ is strictly included in $\mathcal{Z}(L)$ and say that the centralizer is non-trivial.
We will study the structure of $\mathcal{Z}(L)$ as a $\mathbf{C}[L]$-module, as it will allow us to describe $\mathcal{Z}(L)$ by a finite basis. For this purpose we will review next [19, Theorem 1.2]. Let us assume that $\rm ord(L)=n$ and define the $\mathbf{C}[L]$-submodules
@@
If $\mathcal{Z}_{i}$ is non empty then there exists $A_{i}$ in $\mathcal{Z}_{i}$ of minimal order. Observe that $\mathcal{Z}_{0}=\mathbf{C}[L]$ and $A_{0}=1$. By [19, Theorem 1.2], the centralizer $\mathcal{Z}(L)$ is a $\mathbf{C}[L]$-module with basis
@@
The rank of a commutative set of differential operators in $\Sigma[\partial]$ is the greatest common divisor of their orders, see [48] for a discussion about the rank of a set of differential operators. As a consequence of Theorem 2.2, the rank of the centralizer is
@@
Let $L$ be an operator in $\Sigma[\partial]\backslash\mathbf{C}[\partial]$ with non-trivial centralizer $\mathcal{Z}(L)\neq\mathbf{C}[L]$ and consider a basis $\mathcal{B}(L)=\{A_{i},i\in I\}$ of $\mathcal{Z}(L)$ as a $\mathbf{C}[L]$-module. The following are equivalent:
@@
    There exist operators $P$ and $Q$ in $\{L,A_{i}\mid i\in I\}$ of coprime order.
@@ 2.1 Computing ODOs with non-trivial centralizer
Given a differential operator $L$ in $\Sigma[\partial]$ it is an open problem to determine algorithmically if $\mathcal{Z}(L)$ is non-trivial. Although profound and inspiring theoretical studies about the nature of centralizers have been achieved throughout the twentieth century [29], [44],[61], [43],[6], [46], new questions arise to achieve the development of practical algorithms. To generate explicit examples of ODOs with non-trivial centralizer one could use different methods see [63]: Krichever’s explicit formulae for centralizers of rank one with smooth spectral curve; Wilson’s formulae for centralizers of rank one with rational singular spectral curves; the method of deformation of Tyurin parameters from [31], was used by many authors as for instance in [37], [39]; using quintets as in [43], see [46], to obtain differential operators with coefficients in C[[x]]. A new method is proposed in [25] for computable differential extensions $\mathbf{C}\langle\eta\rangle$, where $\eta$ is a transcendental element over $\mathbf{C}\langle\eta\rangle$, say $\eta=\cosh(x)$ or the Weierstrass $\wp$ function.
A crucial reason is that the coefficients of $L$ are solutions of nonlinear differential equations. If we restrict to the case of monic ODOs in normal form (the coefficient of $\partial^{n-1}$ is zero)
@@
In [25], we use Goodearl’s Theorem and our SageMath implementation of GD hierarchies and almost commuting operators in [26] to compute centralizers and generate families of ODOs with non-trivial centralizer.
@@
Consider the order $4$ operator given in  [46, Example 22]. The centralizer of the next operator of order 44
@@ 2.2 In the first Weyl algebra
with $U(x)$, $V(x)$ and $W(x)$ polynomials in $\mathbb{C}[x]$. We proved in [48], Section 6 that given $L_{4}$ irreducible and such that $\mathcal{Z}(L_{4})\not=\mathbb{C}[L_{4}]$, then by [48], Lemma 6.4
$\deg(V)>\max\left\{\frac{1}{2}\deg(U),\frac{1}{2}\deg(W)\right\},$
and by [48], Corollary 6.5
$\mathcal{Z}(L_{4})={\mathbb{C}}[L,B_{2}],$
@@ 3 Spectral curves
By a famous theorem of I. Schur in [54], stated for analytic coefficients, given a monic differential operator $L$ in $\Sigma[\partial]$, if $\rm ord(L)=n$ then the centralizer $\mathcal{Z}((L))$ of $L$ in the ring of pseudo-differential operators is determined by the unique monic $n$-th root $L^{1/n}$ of $L$. For coefficients in a differential ring $R$ this result was proved in [19, Theorem 3.1]
@@
As a consequence, $\mathcal{Z}((L))$ is a commutative ring. To review the long history of this theorem, see  [19, Sections 3 and 4]. By  [19, Lemma 1.1] $\mathcal{Z}(L)$ is a domain and by  [19, Corollary 4.2] it is a commutative domain since
@@
By (7) and (8) the quotient field of $\mathcal{Z}(L)$ is a function field in one variable, whose transcendence degree is one. Moreover any subalgebra $\mathcal{A}$ of $\mathcal{Z}(L)$ has Krull dimension one, it is the affine ring of an algebraic curve ${\rm Spec}(\mathcal{A})$. This fact was first proved by Burchnall and Chaundy [7, 8], and in greater generality by Krichever [30], followed by Mumford [44], Verdier [58], Mulase [43] and [6], Theorem 1.8 with coefficient in $\mathbb{C}((x))$. In this section, we aim to describe the defining ideal of this abstract curve.
@@ 3.1 Defining ideals of spectral curves
We will next define the Burchnall–Chaundy ideal of a finitely generated commutative $\mathbf{C}$-algebra $\mathcal{A}$ in $\Sigma[\partial]$. Without lost of generality we can assume that
@@
with $\rm ord(L)<\rm ord(G_{i})$ and $G_{i}\notin\mathbf{C}[L]$. Observe that $\mathcal{A}\subseteq\mathcal{Z}(L)$.
@@
We will call the elements of the BC ideal BC-polynomials. The next isomorphism holds
@@
Since we are working with differential operators in $\Sigma[\partial]$, then there are no zero divisors in $\mathcal{A}$. Thus $\mathcal{A}$ is an integral domain and $\texttt{BC}(\mathcal{A})$ is a prime ideal. Moreover $\mathcal{A}$ is the coordinate ring of an algebraic curve whose defining ideal is $\texttt{BC}(\mathcal{A})$.
The spectral curve of $\mathcal{A}$ is the irreducible algebraic variety of the prime ideal $\texttt{BC}(\mathcal{A})$, that is
@@
$\mathbf{C}[\Gamma_{\mathcal{A}}]:=\frac{\mathbf{C}[\lambda,\overline{\mu}]}{\texttt{BC}(\mathcal{A})}.$